OpenStudy (anonymous):

What are the greatest common factors of 12,60,102

OpenStudy (anonymous):

well what are the factors of each? 12 = 2*2*3 60 = 2*2*3*5 102 = 2*3*17 so take the one they have in common, 2*3 = 6

OpenStudy (anonymous):

First, lets factor these. 12:1,2,3,4,6,12 60:1,2,3,4,5,6,10,12,15,20,30,60 102:1,2,3,6,17,34,51,102 The greatest one in common is 6.

OpenStudy (anonymous):

should there be more than one? because if the answer is factorS then you're probably looking for the greatest common prime factors.. which would just be 2 and 3.